Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 mph and B is traveling 48 mph. Train A passes the station at 7:15 a.m. If train B passes the same station at 7:45 a.m.,at what time will train B catch up to train A?
Answer by checkley77(12569)

D=RT
D=40T FOR THE A TRAIN.
D=48(T-.5) FOR THE B TRAIN.
BECAUSE THE DISTANCES ARE EQUAL WE SET THE 2 EQUATIONS EQUAL
40T=48(T-.5)
40T=48T-24
40T-48T=-24
8T=24
T=24/8
T=3 HOURS AFTER TRAIN A LEAVES THE STATION, TRAIN B WILL CATCH TRAIN A.
PROOF:
40*3=48(3-.5)
120=48*2.5
120=120